A weather balloon is designed to expand to a maximum radius of 21 m at its working altitude, where the air pressure is 0.030 atm and the temperature is 200 K. If the balloon is filled at atmospheric pressure and 318 K, what is its radius at liftoff?
Answer #1

Va = (4/3)πR³ = 38792 m³

Vsl = Va*[0.030/1.00]*[318/200] = 1850 m³ = (4/3)πR³

=> R = [(3/4π)*1850]^(1/3)

= 7.62 m